You purchased 500 shares of XYZ stock for $41.80 a share. You have received a total of $820 in dividends and $7,280 in proceeds from selling the shares. What is your capital gains yield on this stock? Select one: a. 3.99% b. 32.38% c. 3.92% d. 34.83% e. 38.76%

A stock currently has $2 Earnings Per Shares (EPS). Analysts estimate EPS may grow at 25% per year over the next five years. The investor thinks the stock may sell for a 15 P/ E ratio in five years. What will the stock trade for in five years (from week three)? Select one: a. $93.50 b. $90.50 c. $91.50 d. $99.00 e. $100

The company borrows money at 4% interest rate and the company pays income taxes at 25% rate. Investors wish to earn 15% on its stock. The company has 1 Million shares of common stock outstanding currently trading at $15 per share. The company also has $5 Million debt (bond) outstanding. What is this company’s weighted average cost of capital? Select one: a. 11% b. 12% c. 13% d. 10% e. 9%

You work as analyst for the CFO for a smaller exchange traded company. The CFO says we are getting ready to do a new five-year bond offering tomorrow (bonds have a five year life). We plan selling 50,000 bonds at $1,000 each for $50 Million total. The bonds have a 7% interest rate with interest paid once per year. Unfortunately, the investment banker just called, and they plan pricing this offering for 8% yield to maturity. The bonds will pay 7% interest rate. However, the market demands we pay 8%. The CFO asks you: we planned getting $50 Million, how much we will now get from the offering net of the 2% fee the investment bankers get on the amount of $ they raise for us? Select one: a. $48,454,673 b. $52,045,737 c. $47,043,430 d. $45,783,569 e. $49,052,630

Best Market is planning on paying annual dividends of $1.20, $1.35, and $1.50 over the next 3 years, respectively. After that, Best plans to pay a constant dividend of $1.75 per share each year. To compute the value of Best's stock today, you should first determine the value of the stock at the end of year _____. 
 Select one: a. 2 b. 4 c. 1 d. 0 e. 3 Please answer asap. if possible, please provide steps
Answer #1

Solution:

500 Shares bought at the price of 41.80.

Total investment =500 * 41.80 = 20900

Total Proceeds from selling = 7280 and dividend = 820

Capital Gain Yield = (Capital gain / Total Investment) * 100 = (7280 / 20900) * 100 = .3483 *100 = 34.83%

Option d is the correct answer